OpenStudy (anonymous):

Rationalize the denominator and simplify. 15sqrt(6)/3sqrt(5)

OpenStudy (jdoe0001):

simplify for a fraction, means "get rid of the pesky radical in the denominator" so, we'd do that by simply using that radical and multiplying the whole fraction by it thus \(\bf \cfrac{15\sqrt{6}}{3{\color{blue}{ \sqrt{5}}}}\cdot \cfrac{{\color{blue}{ \sqrt{5}}}}{{\color{blue}{ \sqrt{5}}}}\implies \cfrac{15\sqrt{6}\cdot \sqrt{5}}{3\sqrt{5}\cdot \sqrt{5}}\implies \cfrac{15\sqrt{6\cdot 5}}{3\sqrt{5^2}}\) surely you can simplify it from there

OpenStudy (anonymous):

thanks